OpenStudy (anonymous):

Find the length of the sides of two square whose areas are 405 sq feet and 80 sq feet. Then, find the difference of their lengths?

OpenStudy (anonymous):

A=s^2

OpenStudy (anonymous):

9sqrt(5) and 4sqrt(5)

OpenStudy (anonymous):

difference is 5sqrt(5)

OpenStudy (anonymous):

area of a square = s^2 405=s^2 s=sqrt405=9sqrt5 80=s^2 s=sqrt80=4sqrt5 difference=9sqrt5-4sqrt5=5sqrt5